### v3

# Decomposition of the Tschirnhausen module for coverings on decomposable $\mathbb{P}^{1}$-bundles

Abstract.
In this note, we show that for a smooth algebraic variety $Y$ and a smooth $m$-section $X$ of the $\mathbb{P}^{1}$-bundle
$f:\mathbb{P}(\mathcal{O}_{Y}\oplus\mathcal{O}_{Y}(E))\longrightarrow Y,$
where $E$ is an effective divisor on $Y$ satisfying $H^{1}(Y,\mathcal{O}_{Y}(kE))=0$ for all $k=1,\ldots,m-1$, the Tschirnhausen module of the induced covering $f|_{X}:X\longrightarrow Y$ is completely decomposable. We then apply it to coverings of curves arising in such a way.
The first author was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT) (RS-2024-00352592). The second author was supported by Grant KP-06-N 62/5 of Bulgarian National Science Fund. The third author was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT) (2022R1A2C1005977).
Mathematics Subject Classification (2020): 14E20, 14H30, 14J26, 14H50.
Keywords: Tschirnhausen module, coverings of algebraic varieties, coverings of curves.
1. Introduction and statement
Let $\varphi:X\to Y$ be a finite morphism of degree $m\geq 2$, where $X$ and $Y$ are smooth projective algebraic varieties of dimension $n\geq 1$. Such a covering induces the following short exact sequence of vector bundles on $Y$:
$0\to\mathcal{O}_{Y}\xrightarrow{\varphi^{\sharp}}\varphi_{\ast}\mathcal{O}_{X}\to\mathcal{E}^{\vee}\to 0\,$,
where $\mathcal{E}^{\vee}$ is known as the Tschirnhausen module associated with the cover $\varphi$. The theory of $m:1$ covers in algebraic geometry was developed by Miranda [Mir85] in the case $m=3$, and by Casnati and Ekedahl in general, see [CE96]. Partial results, examples, and applications have since been obtained by many others, including [Kan13], which described the resolution of $\mathcal{O}_{X}$ for $m=3,4,5$ in the case of curves, and [DP22], which showed that every vector bundle on a smooth projective curve arises, up to twist, as a Tschirnhausen module.
In this work, we focus on the situation where the covering arises from a decomposable $\mathbb{P}^{1}$-bundle over the base variety. The main result of the paper is the following theorem.
Theorem 1.
Let $Y$ be a smooth projective variety of dimension $n\geq 1$, and let $E$ and $\Delta$ be effective divisors on $Y$ such that
$H^{1}(Y,\mathcal{O}_{Y}(kE+\Delta))=0\quad\text{ for }k=1,2,\ldots,m-1,$
where $m\geq 2$. Consider the projective bundle $\mathcal{B}:=\mathbb{P}(\mathcal{O}_{Y}\oplus\mathcal{O}_{Y}(E)),$ following Hartshorne’s convention, with projection $f:\mathcal{B}\to Y.$ Let $H$ denote the tautological divisor on $\mathcal{B}$, so that $\mathcal{O}_{\mathcal{B}}(H)\cong\mathcal{O}_{\mathcal{B}}(1)$, and observe that $f_{\ast}\mathcal{O}_{\mathcal{B}}(1)\;\cong\;\mathcal{O}_{Y}\oplus\mathcal{O}_{Y}(E).$ Suppose that
$X\;\in\;\big|mH+f^{\ast}\Delta\big|$
is a smooth, irreducible divisor (an $m$-section of $f$), and let $\varphi:=f|_{X}:X\longrightarrow Y$ be the restriction of the projection $f$ to $X$. Then:
  1. (a)
    The pushforward of the structure sheaf of $X$ decomposes as
    $\varphi_{\ast}\mathcal{O}_{X}\;\cong\;\mathcal{O}_{Y}\;\oplus\;\mathcal{O}_{Y}(-E-\Delta)\;\oplus\;\cdots\;\oplus\;\mathcal{O}_{Y}\bigl(-(m-1)E-\Delta\bigr).$
    In particular, the Tschirnhausen module $\mathcal{E}^{\vee}$ associated to the covering $\varphi:X\to Y$ satisfies
    $\mathcal{E}^{\vee}\;\cong\;\left[\mathcal{O}_{Y}\;\oplus\;\mathcal{O}_{Y}(-E)\;\oplus\;\cdots\;\oplus\;\mathcal{O}_{Y}(-(m-2)E)\right]\otimes\mathcal{O}_{Y}(-E-\Delta).$
  2. (b)
    For the twisted sheaf $\mathcal{O}_{X}\bigl((H-f^{\ast}E)|_{X}\bigr)$, one has
    $\varphi_{\ast}\mathcal{O}_{X}\!\big((H-f^{\ast}E)|_{X}\big)\;\cong\;\mathcal{O}_{Y}\;\oplus\;\mathcal{O}_{Y}(-E)\;\oplus\;\mathcal{O}_{Y}(-2E-\Delta)\;\oplus\;\cdots\;\oplus\;\mathcal{O}_{Y}\bigl(-(m-1)E-\Delta\bigr).$
We discovered this result while studying the construction of components of the Hilbert scheme of curves via $m:1$ covers. The case $m=3$ with $\Delta=0$, was previously obtained by Fujita in [Fuj88]. In our earlier work [CIK24], we considered the case $m=3$, $\operatorname{dim}Y=1$, and $\Delta$ equal to a point, motivated by the need to determine the dimension of the space of first-order deformations of $X$. Subsequently, we realized that the decomposition extends to the more general setting described in Theorem 1. This result can be applied in the construction of components of the Hilbert scheme of curves, as detailed in Theorem $A_{m}$ and Theorem $B_{m}$ of [CIK25].
The proof of Theorem 1 is given in Section 2, while in Section 3 we derive several consequences, some of which have applications to the study of the Hilbert scheme of curves, as can be seen, for instance, in [CIK25].
Conventions and notation
Throughout the paper we work over the field $\mathbb{C}$. By smooth projective variety we mean a smooth, integral, projective scheme of finite type over $\mathbb{C}$. For a vector bundle $\mathcal{V}$ on a variety $Y$, we adopt Hartshorne’s convention for the projectivization $\mathbb{P}(\mathcal{V})$. For additional definitions and results not introduced explicitly in the paper, we refer the reader to [Har77].
2. Proof of Theorem 1
The main tool in the proof is the following proposition.
Proposition 2.
Let $Y$ be a smooth projective variety of dimension $n$, and let $\mathcal{E}$ be a vector bundle of rank $r+1$ on $Y$. Consider the projective bundle
$\mathcal{B}:=\mathbb{P}(\mathcal{E})$
in the sense of Hartshorne’s convention, so that $\mathcal{B}$ is a smooth variety of dimension $n+r$, and the fibers of the projection
$f:\mathcal{B}\to Y$
are projective spaces of dimension $r$. Let $H$ denote the tautological divisor on $\mathcal{B}$, i.e., $\mathcal{O}_{\mathcal{B}}(H)\cong\mathcal{O}_{\mathcal{B}}(1)$, with $f_{\ast}\mathcal{O}_{\mathcal{B}}(H)\cong\mathcal{E}$. Then:
  1. (a)
    The Picard group of $\mathcal{B}$ is given by
    $\operatorname{\operatorname{Pic}}(\mathcal{B})\cong\mathbb{Z}[\mathcal{O}_{\mathcal{B}}(H)]\oplus f^{\ast}\operatorname{\operatorname{Pic}}(Y).$
  2. (b)
    The higher direct images of $\mathcal{O}_{\mathcal{B}}(k):=\mathcal{O}_{\mathcal{B}}(kH)$ satisfy:
    1. (i)
      For all $k\in\mathbb{Z}$,
      $f_{\ast}\mathcal{O}_{\mathcal{B}}(k)\cong\begin{cases}\operatorname{\operatorname{Sym}}^{k}\mathcal{E}&\text{ if }k\geq 0,\\ 0&\text{if }k<0.\end{cases}$
    2. (ii)
      For all $k\in\mathbb{Z}$ and $0<i<r$,
      $R^{i}f_{\ast}\mathcal{O}_{\mathcal{B}}(k)=0.$
    3. (iii)
      For $i=r$,
      $R^{r}f_{\ast}\mathcal{O}_{\mathcal{B}}(k)\cong\begin{cases}0&\text{if }k>-r-1,\\ \bigl(f_{\ast}\mathcal{O}_{\mathcal{B}}(-k-r-1)\bigr)^{\vee}\otimes\det\mathcal{E}^{\vee}&\text{ if }k\leq-r-1.\end{cases}$
Proof.
See [Har77, Exercise 8.4, p. 252]. ∎
We also recall the projection formula [Har77, Exercise 8.3, p. 252], which will be applied repeatedly throughout the proof, without explicit reference. Specifically, if $f:M\to N$ is a morphism of ringed spaces, $\mathcal{F}$ is an $\mathcal{O}_{M}$-module, and $\mathcal{V}$ is a locally free $\mathcal{O}_{N}$-module of finite rank, then
$R^{i}f_{\ast}(\mathcal{F}\otimes f^{\ast}\mathcal{V})\;\cong\;R^{i}f_{\ast}(\mathcal{F})\otimes\mathcal{V}.$
We are now ready to proceed with the proof of the theorem.
Proof of Theorem 1.
First, we give the proof of (a).
Let $X\in|mH+f^{\ast}\Delta|$ be a smooth variety. Apply $f_{\ast}$ to the exact sequence
$0\to\mathcal{O}_{\mathcal{B}}(-mH-f^{\ast}\Delta)\to\mathcal{O}_{\mathcal{B}}\to\mathcal{O}_{X}\to 0.$
Since $m\geq 1$ and $\Delta$ is effective, it follows from Proposition 2 that
$f_{\ast}\mathcal{O}_{\mathcal{B}}(-mH-f^{\ast}\Delta)=0,\quad f_{\ast}\mathcal{O}_{\mathcal{B}}=\mathcal{O}_{Y},\quad\mbox{ and }\quad R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}=0.$
Thus, we obtain the exact sequence
(1) $0\to\mathcal{O}_{Y}\to\varphi_{\ast}\mathcal{O}_{X}\to R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}(-mH-f^{\ast}\Delta)\to 0.$
Applying Proposition 2 again, we find
$R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}(-mH-f^{\ast}\Delta)$ $\cong R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}(-mH)\otimes\mathcal{O}_{Y}(-\Delta)$
$\cong\bigl(f_{\ast}\mathcal{O}_{\mathcal{B}}(m-2)\bigr)^{\vee}\otimes\det\bigl(\mathcal{O}_{Y}\oplus\mathcal{O}_{Y}(E)\bigr)^{\vee}\otimes\mathcal{O}_{Y}(-\Delta)$
$\cong\bigl(\operatorname{Sym}^{m-2}(\mathcal{O}_{Y}\oplus\mathcal{O}_{Y}(E))\bigr)^{\vee}\otimes\mathcal{O}_{Y}(-E-\Delta)$
$\cong\mathcal{O}_{Y}(-E-\Delta)\oplus\cdots\oplus\mathcal{O}_{Y}\bigl(-(m-1)E-\Delta\bigr)\,.$
Since, by assumption,
$h^{1}\bigl(Y,\mathcal{O}_{Y}(kE+\Delta)\bigr)=0\quad\mbox{ for }\quad k=1,\ldots,m-1,$
we have
$\operatorname{Ext}^{1}\bigl(R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}(-mH-f^{\ast}\Delta),\mathcal{O}_{Y}\bigr)\cong H^{1}\bigl(Y,\mathcal{O}_{Y}(E+\Delta)\oplus\cdots\oplus\mathcal{O}_{Y}((m-1)E+\Delta)\bigr)=0.$
Therefore, the exact sequence (1) splits, and we conclude that
$\varphi_{\ast}\mathcal{O}_{X}\cong\mathcal{O}_{Y}\oplus\mathcal{O}_{Y}(-E-\Delta)\oplus\cdots\oplus\mathcal{O}_{Y}\bigl(-(m-1)E-\Delta\bigr).$
The proof of (b) proceeds in a similar way.
Consider the exact sequence
$0\to\mathcal{O}_{\mathcal{B}}\bigl(-(m-1)H-f^{\ast}(E+\Delta)\bigr)\to\mathcal{O}_{\mathcal{B}}(H-f^{\ast}E)\to\mathcal{O}_{X}\bigl((H-f^{\ast}E)_{|_{X}}\bigr)\to 0,$
apply $f_{\ast}$. Since $m\geq 2$, Proposition 2 yields:
  • $\circ$
    $f_{\ast}\mathcal{O}_{\mathcal{B}}\bigl(-(m-1)H-f^{\ast}(E+\Delta)\bigr)\cong f_{\ast}\mathcal{O}_{\mathcal{B}}\bigl(-(m-1)H\bigr)\otimes\mathcal{O}_{Y}(-E-\Delta)=0$,
  • $\circ$
    $f_{\ast}\mathcal{O}_{\mathcal{B}}(H-f^{\ast}E)\cong f_{\ast}\mathcal{O}_{\mathcal{B}}(H)\otimes\mathcal{O}_{Y}(-E)\cong\mathcal{O}_{Y}(-E)\oplus\mathcal{O}_{Y}$, and
  • $\circ$
    $R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}(H-f^{\ast}E)\cong R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}(1)\otimes\mathcal{O}_{Y}(-E)=0$.
Thus, we obtain the exact sequence
(2) $0\to\mathcal{O}_{Y}(-E)\oplus\mathcal{O}_{Y}\to\varphi_{\ast}\mathcal{O}_{X}\bigl((H-f^{\ast}E)_{|_{X}}\bigr)\to R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}\bigl(-(m-1)H-f^{\ast}(E+\Delta)\bigr)\to 0.$
For the term $R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}\bigl(-(m-1)H-f^{\ast}(E+\Delta)\bigr)$, we find:
$R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}\bigl(-(m-1)H-f^{\ast}(E+\Delta)\bigr)$ $\cong R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}\bigl(-(m-1)H\bigr)\otimes\mathcal{O}_{Y}(-(E+\Delta))$
$\cong\bigl(\operatorname{Sym}^{m-3}(\mathcal{O}_{Y}\oplus\mathcal{O}_{Y}(E))\bigr)^{\vee}\otimes\mathcal{O}_{Y}(-(2E+\Delta))$
$\cong\mathcal{O}_{Y}(-2E-\Delta)\oplus\cdots\oplus\mathcal{O}_{Y}(-(m-1)E-\Delta).$
Computing $\operatorname{Ext}^{1}\bigl(R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}(-(m-1)H-f^{\ast}(E+\Delta)),\mathcal{O}_{Y}(-E)\oplus\mathcal{O}_{Y}\bigr)$, we obtain
$\operatorname{Ext}^{1}$ $\bigl(R^{1}f_{\ast}\mathcal{O}_{\mathcal{B}}(-(m-1)H-f^{\ast}(E+\Delta)),\mathcal{O}_{Y}(-E)\oplus\mathcal{O}_{Y}\bigr)$
$\cong H^{1}\bigl(Y,\left[\mathcal{O}_{Y}(2E+\Delta)\oplus\cdots\oplus\mathcal{O}_{Y}((m-1)E+\Delta)\right]\otimes\left[\mathcal{O}_{Y}(-E)\oplus\mathcal{O}_{Y}\right]\bigr)$
$\cong H^{1}\bigl(Y,\left[\mathcal{O}_{Y}(E)\oplus\oplus^{2}_{1}\mathcal{O}_{Y}(2E)\oplus\cdots\oplus\oplus^{2}_{1}\mathcal{O}_{Y}((m-2)E)\oplus\mathcal{O}_{Y}\bigl((m-1)E\bigr)\right]\otimes\mathcal{O}_{Y}(\Delta)\bigr)$
$=0,$
where the vanishing follows from the assumptions of the theorem. Therefore, the exact sequence (2) splits, yielding the desired decomposition.
This completes the proof of the theorem. ∎
3. The case of curves
In this section, we specialize to the case where $Y$ is a smooth projective curve and $E$ is an effective, nonspecial divisor of degree $e$ on $Y$. In this situation, the bundle
$\mathcal{B}=\mathbb{P}\bigl(\mathcal{O}_{Y}\oplus\mathcal{O}_{Y}(E)\bigr)$
is a decomposable ruled surface. Before deriving the corollaries of Theorem 1 in this setting, we recall some classical facts about decomposable ruled surfaces.
The surface $\mathcal{B}$ contains a section $\sigma_{0}$ corresponding to
$\sigma_{0}:\quad\mathcal{O}_{Y}\oplus\mathcal{O}_{Y}(E)\twoheadrightarrow\mathcal{O}_{Y}.$
Let $Y_{0}:=\sigma_{0}(Y)$. Then $Y_{0}$ is the section of minimal self-intersection, satisfying $Y_{0}^{2}=-e$. The surface $\mathcal{B}$ also contains another section $\sigma_{1}$, associated with the surjection
$\sigma_{1}:\quad\mathcal{O}_{Y}\oplus\mathcal{O}_{Y}(E)\twoheadrightarrow\mathcal{O}_{Y}(E).$
If we denote $Y_{1}:=\sigma_{1}(Y)$, then
$Y_{1}\sim Y_{0}+f^{\ast}E\sim H,$
where $H$ is the tautological divisor as in the general case. The intersection numbers are given by
$Y_{0}\cdot H=0,\qquad H^{2}=e.$
Proposition 3.
Suppose that $Y$ is a smooth projective curve of genus $\gamma$, $E$ is a very ample nonspecial divisor of degree $e$, and $\mathcal{B}$, $Y_{0}$, $Y_{1}$, and $H$ are as above.
  1. (i)
    The linear series $|\mathcal{O}_{\mathcal{B}}(H)|$ is base-point-free and defines a morphism
    $\Psi:=\Psi_{|\mathcal{O}_{\mathcal{B}}(H)|}:\mathcal{B}\longrightarrow\mathbb{P}^{R},$
    where $R=e-\gamma+1$.
  2. (ii)
    The morphism $\Psi$ is an isomorphism away from $Y_{0}$, which it contracts to a point.
  3. (iii)
    Geometrically, the image $F:=\Psi(\mathcal{B})$ is a cone in $\mathbb{P}^{R}$ over a smooth curve $Y_{e}\cong Y$ of degree $e$, embedded in $\mathbb{P}^{R-1}$. The lines in the ruling of $F$ are the images of the fibers $f^{-1}(q)$ for $q\in Y$.
Proof.
The proof is obtained easily using [Har77, Ex. V.2.11, p.385] and [FP05, Proposition 23]. For additional details the reader can refer to [CIK25, Lemma 2.2]. ∎
In the above setting, let $P$ denote the vertex of the cone $F\subset\mathbb{P}^{R}$. We also note that in Corollary 4 and Corollary 6 that follow, it is assumed that the divisor $E$ on the curve $Y$ of genus $\gamma$ is nonspecial and very ample.
Corollary 4.
Let $X\subset\mathcal{B}$ be an irreducible smooth curve in the linear series $|mH|$, and let $X_{me}:=\Psi(X)\subset F$ be the image of $X$ under $\Psi$. Then:
  1. (a)
    $X_{me}$ is a smooth curve of degree $me$ and genus
    $g=\binom{m}{2}e+m\gamma+1-m,$
    cut from $F$ by a degree $m$ hypersurface in $\mathbb{P}^{R}$;
  2. (b)
    if $\phi:X_{me}\to Y_{e}$ is the morphism induced by projection from $P$ to the hyperplane in $\mathbb{P}^{R}$ containing $Y_{e}$, then
    (3) $\phi_{\ast}\mathcal{O}_{X_{me}}\;\cong\;\mathcal{O}_{Y_{e}}\oplus\mathcal{O}_{Y_{e}}(-1)\oplus\cdots\oplus\mathcal{O}_{Y_{e}}\bigl(-(m-1)\bigr).$
Proof.
Since $H\cdot Y_{0}=0$, the morphism $\Psi$ maps each smooth curve in the linear system $|kH|$ on $\mathcal{B}$ (for $k\geq 1$) isomorphically onto a curve on $F$, cut out by a hypersurface of degree $k$ in $\mathbb{P}^{R}$. In particular, $Y_{e}\cong Y$ appears as a hyperplane section of $F$, and $X_{me}\cong X$ as a hypersurface section of degree $m$. The formula for the genus then follows directly from the adjunction formula. Also, $\deg X_{me}=mH\cdot H=me$. This establishes (a).
For part (b), note that the morphism $\phi:X_{me}\to Y_{e}$ corresponds, via the diagram below, to the morphism $\varphi:X\to Y$ from Theorem 1 with $\Delta=0$:
[…]
Moreover,
$\bigl(\Psi_{|_{Y_{1}}}\bigr)^{\ast}\mathcal{O}_{Y_{e}}(1)\;\cong\;\mathcal{O}_{Y_{1}}\bigl(H_{|_{Y_{1}}}\bigr)\;\cong\;\bigl(f_{|_{Y_{1}}}\bigr)^{\ast}\mathcal{O}_{Y}(E).$
Thus, applying part (a) of Theorem 1 yields the decomposition (3), completing the proof. ∎
If $q\in Y$ is a point, then the fiber $f^{\ast}q=f^{-1}(q)$ of $f:\mathcal{B}\to Y$ meets the section $Y_{0}\sim H-f^{\ast}E$ in a unique point, denoted
$q_{0}:=Y_{0}\cap f^{\ast}q.$
If $X\in|mH+f^{\ast}q|$ is a smooth curve, then
(5) $\mathcal{O}_{X}\big((H-f^{\ast}E)|_{X}\big)\;\cong\;\mathcal{O}_{X}\big(Y_{0}|_{X}\big)\;\cong\;\mathcal{O}_{X}(q_{0}).$
Moreover, for any $z\in Y$, the morphism $\Psi$ maps the fiber $f^{\ast}z=f^{-1}(z)$ onto a line in the ruling of $F$.
Remark 5.
As noted in [CG99, p. 226], the linear system $|mH+f^{\ast}q|$ contains a smooth element provided there exist reduced divisors $\tilde{E}\in|E|$ and $G\in|mE+q|$ such that $q\in\tilde{E}$ and $\tilde{E}\cap G=\varnothing$. Since $E$ is a nonspecial divisor on $Y$, the system $|mE+q|$ is base-point free, so $q$ is not a fixed point. When $E$ is nonspecial and $h^{0}\bigl(Y,\mathcal{O}_{Y}(E)\bigr)\geq 2$, Bertini-type arguments guarantee that one can choose $\tilde{E}\sim E$ with $q\in\tilde{E}$ and a reduced divisor $G\sim mE+q$ disjoint from $\tilde{E}$. Hence, whenever $E$ is a nonspecial divisor on $Y$ with $h^{0}\bigl(Y,\mathcal{O}_{Y}(E)\bigr)\geq 2$, the general element of $|mH+f^{\ast}q|$ is smooth and irreducible.
Corollary 6.
Let $\Delta=q$ be a point on $Y$, and let $X\subset\mathcal{B}$ be a smooth curve in the linear series $|mH+f^{\ast}q|$. Denote by $X_{me+1}:=\Psi(X)\subset F$ the image of $X$ under $\Psi$. Let $l_{q}$ be the line in the ruling of $F$ corresponding to the image of $f^{\ast}q$, and let $Q:=l_{q}\cap Y_{e}$. Then:
  1. (a)
    $X_{me+1}$ is a smooth curve of degree $me+1$ and genus
    $g=\binom{m}{2}e+m\gamma,$
    which passes through the vertex $P$ of $F$ with multiplicity one;
  2. (b)
    if $\phi:X_{me+1}\to Y_{e}$ is the morphism induced by projection from $P$ onto the hyperplane in $\mathbb{P}^{R}$ containing $Y_{e}$, then:
    • $\bullet$
      (6) $\phi_{\ast}\mathcal{O}_{X_{me+1}}\;\cong\;\mathcal{O}_{Y_{e}}\oplus\Bigl(\mathcal{O}_{Y_{e}}(-1)\oplus\cdots\oplus\mathcal{O}_{Y_{e}}\bigl(-(m-1)\bigr)\Bigr)\otimes\mathcal{O}_{Y_{e}}(-Q),$
    • $\bullet$
      (7) $\phi_{\ast}\mathcal{O}_{X_{me+1}}(P)\;\cong\;\mathcal{O}_{Y_{e}}\oplus\mathcal{O}_{Y_{e}}(-1)\oplus\Bigl(\mathcal{O}_{Y_{e}}(-2)\oplus\cdots\oplus\mathcal{O}_{Y_{e}}\bigl(-(m-1)\bigr)\Bigr)\otimes\mathcal{O}_{Y_{e}}(-Q).$
Proof.
Part (a) follows by arguments analogous to those used in Corollary 4. For the statement about the multiplicity at $P$, see also [CIK24, Prop. 3 and 4].
For part (b), note that the projection from $P$ to the hyperplane containing $Y_{e}$ induces the morphism $\phi:X_{me+1}\to Y_{e}$. This corresponds to the morphism $\varphi:X\to Y$ from Theorem 1, via a commutative diagram analogous to (4). The claimed decompositions (6)–(7) then follow from (5) and Theorem 1. ∎
The previous corollaries allow us to derive the following result.
Proposition 7.
Let $Y_{e}\subset\mathbb{P}^{R-1}$ be a smooth curve of genus $\gamma\geq 2$ and degree $e\geq 2\gamma-1$, with $R=e-\gamma+1$. Let $F\subset\mathbb{P}^{R}$ be the non-degenerate cone over $Y_{e}$ with vertex $P$. Fix an integer $m\geq 2$, and let $X_{d}\subset F$ be a smooth curve of degree $d$ and genus $g$. Denote by
$\phi:X_{d}\longrightarrow Y_{e}$
the morphism induced by projection from $P$ onto the hyperplane containing $Y_{e}$. Then:
  1. (a)
    If $d=me$ and $g=\binom{m}{2}e+m\gamma+1-m$, then $X_{d}$ is cut out on $F$ by a hypersurface of degree $m$. Moreover, the direct image $\phi_{\ast}\mathcal{O}_{X_{d}}$ decomposes as in (3).
  2. (b)
    If $d=me+1$ and $g=\binom{m}{2}e+m\gamma$, then $X_{d}$ is algebraically equivalent to the intersection of $F$ with a hypersurface of degree $m$, together with a line from the ruling of $F$. The tangent line $\tau_{P}$ to $X_{d}$ at $P$ belongs to the ruling of $F$ and meets $X_{d}$ at $P$ with multiplicity two. Setting $Q:=\tau_{P}\cap Y_{e}$, the pushforwards $\phi_{\ast}\mathcal{O}_{X_{d}}$ and $\phi_{\ast}\mathcal{O}_{X_{d}}(P)$ satisfy (6) and (7), respectively.
Proof.
Let $S$ be the blow-up of $F$ at its vertex $P$. Then $S$ can be identified with the decomposable ruled surface
$\mathbb{P}\bigl(\mathcal{O}_{Y_{e}}\oplus\mathcal{O}_{Y_{e}}(1)\bigr)$
over $Y_{e}$. Let $f:S\to Y_{e}$ be the natural projection, and let $H$ denote the tautological divisor on $S$, so that
$f_{\ast}\mathcal{O}_{S}(H)\;\cong\;\mathcal{O}_{Y_{e}}\oplus\mathcal{O}_{Y_{e}}(1).$
Since $F$ is the image of $S$ under the morphism $\Psi$ associated with the linear system $|H|$, it suffices to show that:
  • in case (a), the curve $X_{d}$ is the image of a smooth curve $X\sim mH$ on $S$;
  • in case (b), the curve $X_{d}$ is the image of a smooth curve $X\sim mH+f^{\ast}Q$ on $S$.
We treat case (a); case (b) was established in [CIK24, Proposition 5].
Assume $d=me$ and $g=\binom{m}{2}e+m\gamma+1-m$ as in (a). Let $X\subset S$ be the strict transform of $X_{d}$, and let $Y_{0}:=\Psi^{-1}(P)$ denote the section of minimal self-intersection on $S$. Then $Y_{0}^{2}=-e$, and
$H\;\sim\;Y_{0}+f^{\ast}\xi,$
where $\xi$ is the hyperplane section divisor on $Y_{e}$, i.e. $\mathcal{O}_{Y_{e}}(\xi)\cong\mathcal{O}_{Y_{e}}(1)$.
Since $S$ is generated by $Y_{0}$ and pullbacks of divisors from $Y_{e}$, we may write
$X\;\sim\;aY_{0}+f^{\ast}\beta$
for some divisor $\beta$ on $Y_{e}$. From the degree condition
$H\cdot X=\deg X_{d}=me,$
we obtain $\deg\beta=me$.
Applying the adjunction formula gives
$\bigl(-2Y_{0}+f^{\ast}\omega_{Y_{e}}-f^{\ast}\xi+aY_{0}+f^{\ast}\beta\bigr)\cdot\bigl(aY_{0}+f^{\ast}\beta\bigr)=2g-2=m(m-1)e+2m(\gamma-1),$
which simplifies to
(8) $a(a-2)e+me(a-2)+a(2\gamma-2-e+me)\;=\;m\bigl((m-1)e+2(\gamma-1)\bigr).$
Solving (8) for $a$, we find the roots
$a=m\quad\text{or}\quad a=m+1+\tfrac{2(\gamma-1)}{e}.$
Since $a$ must be integer and $e\geq 2\gamma-1$ by hypothesis, the second solution is excluded. Thus $a=m$, and hence
$X\;\sim\;mY_{0}+f^{\ast}\beta.$
Next, observe
$X\cdot Y_{0}=(mY_{0}+f^{\ast}\beta)\cdot Y_{0}=-me+\deg\beta=0,$
so the curves $X$ and $Y_{0}$ are disjoint. Consequently,
$(f^{\ast}\beta)_{|Y_{0}}\;\sim\;-m(Y_{0}|_{Y_{0}}).$
Identifying $Y_{0}\cong Y_{e}$ via the isomorphism $j:={f}_{|_{Y_{0}}}:Y_{0}\to Y_{e}$, we have
$Y_{0}|_{Y_{0}}\;\sim\;-j^{\ast}\xi\quad\text{and}\quad(f^{\ast}\beta)_{|Y_{0}}\;\sim\;j^{\ast}\beta.$
It follows that $j^{\ast}\beta\sim mj^{\ast}\xi$, hence $\beta\sim m\xi$ on $Y_{e}$. Therefore,
$X\;\sim\;m(Y_{0}+f^{\ast}\xi)\;\sim\;mH.$
The claim in (a) now follows from Corollary 4. ∎
Corollary 8.
In the setting of Proposition 7, let $X_{d}\subset F$ be a smooth curve of degree $d$ such that the morphism $\phi:X_{d}\to Y_{e}$ is an m:1m\!:\!1 cover. Then:
  1. (a)
    If $X_{d}$ does not pass through $P$, then $d=me$ and $g=\binom{m}{2}e+m\gamma+1-m$. In particular, case (a) of Proposition 7 applies.
  2. (b)
    If $X_{d}$ passes through $P$, then $d=me+1$ and $g=\binom{m}{2}e+m\gamma$. In particular, case (b) of Proposition 7 applies.
Proof.
Let $S$, $f$, $X$, $H$, $Y_{0}$, $\xi$, and $\Psi$ be as in the proof of Proposition 7. Write
$X\sim aY_{0}+f^{\ast}\beta,$
for some integer $a$ and some divisor $\beta$ on $Y_{e}$. Since $\Psi$ maps fibers $f^{\ast}z$ (for $z\in Y_{e}$) to lines in the ruling of $F$, and $\phi$ is the projection from $P$ onto the hyperplane containing $Y_{e}$, the assumption that $\phi$ is an m:1m\!:\!1 cover implies
$m=X\cdot f^{\ast}z=a.$
(a) If $X_{d}$ does not contain $P$, then
$Y_{0}\cdot(mY_{0}+f^{\ast}\beta)=0,$
hence $Y_{0}\cdot f^{\ast}\beta=me$. By the adjunction formula we obtain
$2g-2=\bigl(-2Y_{0}+f^{\ast}\omega_{Y_{e}}-f^{\ast}\xi+mY_{0}+f^{\ast}\beta\bigr)\cdot(mY_{0}+f^{\ast}\beta)=m(m-1)e+2m(\gamma-1).$
Thus $g=\binom{m}{2}e+m\gamma+1-m$. The degree $d$ of $X_{d}$ is
$d=X\cdot H=(mY_{0}+f^{\ast}\beta)\cdot(Y_{0}+f^{\ast}\xi)=me.$
(b) If $X_{d}$ contains $P$, then
$Y_{0}\cdot(mY_{0}+f^{\ast}\beta)=1,$
since $\Psi$ contracts $Y_{0}$ to $P$. Hence $Y_{0}\cdot f^{\ast}\beta=me+1$. By adjunction,
$2g-2=m(m-1)e+2m\gamma-2,$
so $g=\binom{m}{2}e+m\gamma$. The degree $d$ of $X_{d}$ is
$d=(mY_{0}+f^{\ast}\beta)\cdot(Y_{0}+f^{\ast}\xi)=me+1.$
Finally, we include one more statement, concerning plane curves, which is derived easily using similar arguments as above.
Proposition 9.
Let $X_{m}\subset\mathbb{P}^{2}$ be a smooth plane curve of degree $m\geq 2$, let $L\cong\mathbb{P}^{1}$ be a line in the plane, and let $P\in\mathbb{P}^{2}\setminus L$ be a point not lying on $L$. Consider the morphism
$\phi:X_{m}\longrightarrow L$
induced by projection from $P$ onto $L$. Then:
  1. (a)
    If $P\notin X_{m}$, then $\phi$ is an m:1m\!:\!1 covering, and
    $\phi_{\ast}\mathcal{O}_{X_{m}}\;\cong\;\mathcal{O}_{\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}}(-1)\oplus\cdots\oplus\mathcal{O}_{\mathbb{P}^{1}}\bigl(-(m-1)\bigr),$
    where we identify $\mathcal{O}_{L}$ with $\mathcal{O}_{\mathbb{P}^{1}}$.
  2. (b)
    If $P\in X_{m}$, $m\geq 3$, and the tangent line to $X_{m}$ at $P$ meets $X_{m}$ with multiplicity two, then $\phi$ is an (m1):1(m-1)\!:\!1 covering, and:
    • $\bullet$
      $\phi_{\ast}\mathcal{O}_{X_{m}}\;\cong\;\mathcal{O}_{\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{1}}(-3)\oplus\cdots\oplus\mathcal{O}_{\mathbb{P}^{1}}\bigl(-(m-1)\bigr),$
    • $\bullet$
      $\phi_{\ast}\mathcal{O}_{X_{m}}(P)\;\cong\;\mathcal{O}_{\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{1}}(-3)\oplus\cdots\oplus\mathcal{O}_{\mathbb{P}^{1}}\bigl(-(m-1)\bigr).$
Proof.
Let $S$ be the blow-up of $\mathbb{P}^{2}$ at the point $P$. Then $S\cong\mathbb{F}_{1}=\mathbb{P}\bigl(\mathcal{O}_{\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}}(1)\bigr)$, the first Hirzebruch surface, which can be viewed as the projectivization $S\cong\mathbb{P}\bigl(\mathcal{O}_{L}\oplus\mathcal{O}_{L}(1)\bigr)$ with projection map $f:S\to L$. Let $H$ denote the tautological divisor on $S$, and let $X\subset S$ be the strict (proper) transform of $X_{m}$. Then:
  • in case (a) we have $X\sim mH$,
  • in case (b) we have $X\sim(m-1)H+f^{\ast}Q$, where $Q$ is the point where the tangent line to $X_{m}$ at $P$ meets $L$.
The assertions now follow from Corollary 4 and Corollary 6, respectively. ∎
Remark 10.
In all of the above cases, the decomposition of the Tschirnhausen module is of Veronese type, namely
$\mathcal{E}^{\vee}\;\cong\;\bigl(\mathcal{O}_{Y}\oplus L\oplus\cdots\oplus L^{m-2}\bigr)\otimes M$
for some line bundles $L$ and $M$ on $Y$. It would be interesting to find examples of such decompositions where $X$ is not an $m$-section of a decomposable $\mathbb{P}^{1}$-bundle over $Y$.